OpenStudy (anonymous):

Trig help.. Simplify: (sin Θ − cos Θ)2 + (sin Θ + cos Θ)2 1 2 sin2 Θ cos2 Θ

OpenStudy (anonymous):

Where it says "2" they are raised to the power

OpenStudy (anonymous):

\[ (\sin Θ − \cos Θ)^2 + (\sin Θ + \cos Θ)^2\]\[ \sin^2 Θ −2\sinΘ\cosΘ+ \cos^2 Θ+ (\sin Θ + \cos Θ)^2\]\[ \sin^2 Θ + \cos^2 Θ−2\sinΘ\cosΘ+ (\sin Θ + \cos Θ)^2 \]\[1−2\sinΘ\cosΘ+ (\sin Θ + \cos Θ)^2 \]\[1−2\sinΘ\cosΘ+ \sin ^2Θ +2\sinΘ\cosΘ+ \cos ^2Θ\]\[1−2\sinΘ\cosΘ+2\sinΘ\cosΘ+ \sin ^2Θ + \cos ^2Θ\]\[1−2\sinΘ\cosΘ+2\sinΘ\cosΘ+ 1\]\[2\]

OpenStudy (anonymous):

hope this helps.

OpenStudy (anonymous):

short and clear.

OpenStudy (anonymous):

You are a life saver.

OpenStudy (anonymous):

perhaps... tnx